OpenStudy (anonymous):

write equation shape of y=squareroot x but shifted left 3 units

OpenStudy (mathmale):

Suppose I gave you the function y=f(x) = x^2 and asked you to graph it. Then suppose I asked you to shift the same shape (graph) 2 units to the left. The proper form for the new function you've created would be g(x)=(x+2)^2. Supposing you were to move the graph of f(x)=x^2 1 unit to the right: Your new function would be h(x)= (x-1)^2. Use this info to answer this question.